12- Your roommate’s long hours in Chem lab finally paid off—she discovered a secret formula that lets people do an hour’s worth of studying in 5 minutes. So far, she’s sold 200 doses, and faces the following average-total cost schedule:

Q Average Total Cost
199 $199
200 200
201 201





If a new customer offers to pay your roommate $300 for one dose, should she make one more? Explain.
1
Expert's answer
2019-05-01T10:34:03-0400

If she sell 200 doses for $300 for one dose, then total profit is: TR = (P - ATC)*Q = (300 - 200)*200 = $20,000.

If she sell 201 doses for $300 for one dose, then total profit is: TR = (300 - 201)*201 = $19,899.

So, she should not make one more, because her profits will decrease.